Please help with this problem:Solve the linear system using elimination.
-3x-y=8
7x+y=-12
thank you

Answer by mananth(16949)

-3 x -1 y = 8 .............1
7 x + 1 y = -12 .............2
Eliminate y
Add the two equations
4 x = -4
/ 4
x = -1
plug value of x in (1)
-3 x + -1 y = 8
3 + -1 y = 8
-1 y = 8 -3
-1 y = 5
y = -5
m.ananth@hotmail.ca